G-Blast Gaming members pay a monthly fee that allows unlimited hours of online gaming. G-Blast hires you to consult on the best membership pricing scheme. You estimate its typical user’s inverse demand for streaming gaming as P = 600 – 20Q, where Q is hours of streaming. The company can supply an additional hour of streaming to another member at zero marginal cost. What is the highest price that G-Blast should charge for a membership fee?
Answer #1

P= 600-20Q

Pm= Price when Q is zero= 600

G-Blast could charge the maximum membership which is equals to the consumer surplus so that all the surplus in the market could be acquired by the G-Blast.

For consumer surplus, find P and Q. For P and Q:

P= MC

600-20Q= 0

600= 20Q

Q*= 600/20= 30

Use Q= 30 in Demand equation:

P*= 600-20Q= 600-600= 0 (Price for additional hour)

Consumer surplus= Area above price and below demand curve

Consumer surplus= (1/2)(Pm-P*)(Q*)= (1/2)(600-0)(30)= 300 x 30= 9000

G-Blast should charge for a membership fee= 9000